OpenStudy (anonymous):

xsq2+6x+8=0 separate the two values with a comma

OpenStudy (solomonzelman):

\[x \sqrt{2}~+~6x~+~8~=~0~~~~~~~~~~~~~~~~~~~this?\]

OpenStudy (solomonzelman):

move \[move~~~~~~"6x+8"~~~~t o~~the~~l e ft~~(by~~subtracting~~6x~~and~~then~~8\]\[from~~both~~sides,~~then)~~squ are~~both~~sides.\]

OpenStudy (solomonzelman):

\[Shortcut:~~~~~x \sqrt{2}~+6x+8-(6x+8)=0-(6x+8)\]\[x \sqrt{2}=6x+8,~~~~~~~~~~~now,~~square~~both~~sides,\]\[x^2 \times 2~=~36x^2+96x+64,~~~~which~~is,~~~~~2x^2=36x^2+96x+64\]\[factor~~out~~of~~2,~~~~~x^2=18x^2+48x+32,~~~~now~~you~~can~~do~~i t.\]